OpenStudy (anonymous):

Express answer in exact form. A segment of a circle has a 120 degree arc and a chord of 8 square root of three in. Find the area of the segment.

OpenStudy (anonymous):

A segment of a circle has a 120 arc and a chord of 8in. Find the area of the segment. \[A=(\frac{ aa}{ b }\pi - \sqrt[aa]{b}\]

OpenStudy (anonymous):

the double a means there are two numbers there

OpenStudy (anonymous):

inches squared at the end of the diagram

terenzreignz (terenzreignz):

A radius would be nice.

OpenStudy (anonymous):

it doesnt give me one

terenzreignz (terenzreignz):

It gives you enough info to get the radius though... The arclength L with a given angle A is given by \[\huge L = \frac{A}{360^o}\times2\pi r\]

OpenStudy (anonymous):

so it would be 120/360*2pi

terenzreignz (terenzreignz):

A = 120 degress... and L = ?

OpenStudy (anonymous):

L= \[8\sqrt{3}\]

terenzreignz (terenzreignz):

good. So.. \[\huge 8\sqrt3=\frac{120^o}{360^o}\times2\pi r\] And now, just solve for r.

OpenStudy (anonymous):

i got 2.1

terenzreignz (terenzreignz):

hmm... best not to approximate yet, keep the pi, keep the radical. Do the approximating at the end, with the final answer.

OpenStudy (anonymous):

im so confused